# Optimisation of the vertex detector and measurement of Higgs decays to second-generation quarks at the CEPC

### v2

@@ 1 Introduction
Measurements of the $H\to c\bar{c}$ and $H\to s\bar{s}$ directly test the Yukawa couplings between the Higgs and the second-generation quarks. The SM predicts the $H\to c\bar{c}/s\bar{s}$ has a branching ratio of $2.9\%$$2.3\times 10^{-4}$ SM_BR; Hss_BR; PDG. Notably, these correspond to about 120120k/11k $H\to c\bar{c}/s\bar{s}$ events at the CEPC when considering all Higgs production modes. With the combination of high-performance detectors and advanced particle reconstruction algorithms, the $H\to c\bar{c}$ is expected to be measured to a percent-level of accuracy, while discovering the $H\to s\bar{s}$ decay mode is a critical and promising possibility at the future Higgs factory.
@@ 2 CEPC Vertex Detector and Samples
The CEPC baseline detector adopts a Particle Flow Oriented design Ruan_ArborPerf_2018, emphasizing the separation and reconstruction of every final state particle, and measures them in the most suited sub-detectors. A crucial component of the CEPC detector is its vertex system. As referenced in the CEPC Conceptual Design Report (CDR) CEPC_CDR_Phy, the vertex system consists of six concentric cylindrical layers of silicon pixel sensors (Figure 1). The first barrel layer is positioned merely 16 mm from the beam line, and with a per-layer material budget of only $0.15\%\,X_{0}$, it attains a spatial resolution of $2.8\,\mu\text{m}$ for high precision flavour tagging. In the subsequent CEPC TDR CEPC2025TDR, the Machine Detector Interface region is explicitly modelled, including a 0.35 $mm$ beryllium beam pipe with a 10 $\mu m$ inner gold coating and 0.2 $mm$ cooling-water, corresponding to about 0.454% $X_{0}$ within the vertex acceptance. Each vertex layer contributes approximately 0.06–0.28% $X_{0}$, dominated by the sensors and local supports. The CDR configuration used in this study thus represents an idealized, low-mass scenario suitable for conceptual optimization of the vertex radius and spatial resolution.
@@ 3.2 Performance of jet origin identification
The JOI based on the ParticleNet architecture Qu:2019gqs is designed to determine the colored particle from which a jet originates. It classifies jets into 11 flavour states ($b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, and gluon) by leveraging kinematic information and particle-species identification. Inputs to the algorithm include track impact parameters, energy/momentum distributions, and particle types (e.g. $e^{+}/e^{-}$, $\mu^{+}/\mu^{-}$, $\pi^{+}/\pi^{-}$, $K^{+}/K^{-}$).
Figure 3 shows two complementary views of the JOI performance. The left panel displays the normalized 11-dimensional jet flavour migration matrix (denoted as $M_{11}$), providing a global overview of jet tagging efficiency and misidentification.
The right panel illustrates the dedicated performance of $b$-jets for optimised analysis, showing the relation between the $b$-tagging efficiency and the mis-identification rates of other jet flavours. Here, the $b$-tagging efficiency denotes the fraction of true $b$-jets correctly identified, while the mis-ID rate corresponds to non-$b$ jets misclassified as $b$-jets. For a $b$-tagging efficiency in the range $0.3$ to $0.8$, the JOI achieves a misidentification rate that is 1-3 orders of magnitude lower than that of the multivariable analysis approach-XGBoost method refTDR_report.
@@
From the migration matrices at varying $R_{\mathrm{rad}}$ (see figure 4), we extract key flavour-tagging metrics as shown on the left of figure 5. As $R_{\mathrm{rad}}$ increases from $0.5$ to $2$, the $b\to c$, $g\to c$ and $c\to s$ misidentification rate rises from about $4\%$ to $7\%$, Meanwhile, the $c\to c$ tagging efficiency drops from $81\%$ to $70\%$, and $s\to s$ from $60\%$ to $54\%$. While the $g\to s$ rate remains stable around $14\%$. The trends indicate that reducing the vertex detector’s innermost radius significantly lowers the $b\to s$ and $c\to s$ misidentification rates, reflecting the crucial role of the vertex in distinguishing heavy-flavour jets from strange jets. In contrast, the impact of vertex detector optimisation on gluon-strange discrimination is limited as gluon jets predominantly originate from the primary vertex, and vertex detector upgrades have little effect on their identification.
@@
To further probe the global performance of JOI, we characterize its behaviour via the trace of $M_{11}$, denoted as $\mathrm{Tr}(M)$. This trace, which quantifies the total correct classification probability across all quark flavour categories, is analysed as a function of the geometry scaling factors $\log_{2}(\text{new}/\text{baseline})$ for both $R_{\rm rad}$ and $R_{\rm res}$. As shown in right of figure  5, a linear regression to the $\mathrm{Tr}(M)$ yields:
@@
3.3 Performance of higgs benchmarks: $H\to c\bar{c}$ and $H\to s\bar{s}$ in $\nu\bar{\nu}H$ channel
@@
with $S_{f\bar{f}}\in[0,1]$ and $\sum_{f}S_{f\bar{f}}=1$. For a given threshold on $S_{f\bar{f}}$, the efficiency $\epsilon$ is defined as the number of true signal events that pass the $S_{ff}$ selection divided by the total number of simulated signal events, and the purity $P$ is defined as the fraction of selected events that truly originate from the target decay channel out of all selected events, including both signal and background contributions. Both quantities are therefore derived consistently from the posterior $S_{f\bar{f}}$.
@@ 4 Summary and Discussion
Our study is based on fast simulations and currently considers only the major backgrounds from other Higgs decay modes, representing an optimistic yet well-motivated scenario. In a realistic environment, additional beam-induced and other Standard Model backgrounds would further dilute the sensitivity. However, studies at CEPC and other $e^{+}e^{-}$ collider designs indicate that these effects are expected to be under control, leading to subdominant contributions to the vertex occupancy and flavour-tagging performance. Therefore, the results presented here can be regarded as upper-limit estimates, corresponding to the performance achievable under ideal background conditions. Comparisons with full (Geant4-based) simulations confirm the reliability of our fast-simulation approach, with discrepancies at the 5% level.
@@ Appendix A Comparison between Delphes and Geant4
Appendix B Beam-induced background
The influence of beam-induced background (BIB) on tracking performance has been qualitatively evaluated in the CEPC TDR CEPC2025TDR. As shown in Fig. 10, a direct comparison of the impact-parameter resolutions ($d_{0}$ and $z_{0}$) with and without background overlay demonstrates no noticeable degradation across the full momentum range. Therefore, the pixel size and spatial resolution variations studied in this work remain valid, and the overall conclusions on vertex detector optimization are not affected by beam background effects.

### v3

@@ 1 Introduction
Measurements of the $H\to c\bar{c}$ and $H\to s\bar{s}$ directly test the Yukawa couplings between the Higgs and the second-generation quarks. The SM predicts the $H\to c\bar{c}/s\bar{s}$ has a branching ratio of $2.9\%$$2.3\times 10^{-4}$ SM_BR; Hss_BR; PDG. Notably, these correspond to about 120k $H\to c\bar{c}$ events and about 1k $H\to s\bar{s}$ events at the CEPC when considering all Higgs production modes. Among these, approximately one quarter are produced in the $\nu\bar{\nu}H$ final state from $ZH\,(Z\to\nu\bar{\nu})$ and $WW$-fusion processes. This channel offers a particularly clean environment in which all visible particles originate from the Higgs decay, and therefore provides a convenient benchmark for the simulation study used in this work. With the combination of high-performance detectors and advanced particle reconstruction algorithms, the $H\to c\bar{c}$ is expected to be measured to a percent-level of accuracy, while discovering the $H\to s\bar{s}$ decay mode is a critical and promising possibility at the future Higgs factory.
@@ 2 CEPC Vertex Detector and Samples
The CEPC baseline detector adopts a Particle Flow Oriented design Ruan_ArborPerf_2018, emphasizing the separation and reconstruction of every final state particle, and measures them in the most suited sub-detectors. A crucial component of the CEPC detector is its vertex system. As referenced in the CEPC Conceptual Design Report (CDR) CEPC_CDR_Phy, the vertex system consists of six concentric cylindrical layers of silicon pixel sensors (Figure 1). The first barrel layer is positioned merely 16 mm from the beam line, and with a per-layer material budget of only $0.15\%\,X_{0}$, it attains a spatial resolution of $2.8\,\mu\text{m}$ for high precision flavour tagging. The beam pipe dominates the upstream material of the vertex detector. In the CEPC TDR design CEPC2025TDR, the Machine Detector Interface(MDI) region is explicitly modelled, including a 0.35 mm beryllium beam pipe with a 10 $\mu m$ inner gold coating and a 0.2 mm cooling-water layer, corresponding in total to about 0.45% $X_{0}$. Each vertex layer contributes approximately 0.06 – 0.28% $X_{0}$, dominated by the sensors and local supports.
The beam-induced background is also highly relevant to the design and operation of the vertex detector, as its hits can dominate the occupancy and bandwidth of the detectors. The CEPC TDR provides a dedicated simulation of the beam-induced background, showing that the impact on the impact-parameter resolution is under control. Further details are provided in Appendix B. Hence, the beam-induced background is neglected in this study.
@@ 3.2 Performance of jet origin identification
The JOI based on the ParticleNet architecture Qu:2019gqs is designed to determine the coloured particle from which a jet originates. For each jet, the algorithm takes the kinematic variables of the reconstructed final-state particles, their particle-type information, the impact-parameter measurements of charged tracks, their energy and momentum distributions, as inputs. It then computes likelihood scores for eleven jet-flavour categories ($b$, $\bar{b}$, $c$, $\bar{c}$, $s$, $\bar{s}$, $u$, $\bar{u}$, $d$, $\bar{d}$, and gluon), which are subsequently employed to construct flavour-tagging observables.
Figure 3 shows the performance of the JOI algorithm. The left-hand panel displays the normalized 11-dimensional jet-flavour migration matrix ($M_{11}$), obtained by assigning each jet to the flavour category with the highest score, thus giving an overview of the tagging and misidentification pattern among all flavours. The right-hand panel provides a complementary flavour-specific view by presenting the $b$-jet tagging performance, obtained from scanning working points with different $b$-jet efficiencies. For each chosen $b$-jet efficiency, the corresponding misidentification rates of the other flavour jets are shown. The $b$-tagging efficiency is defined as the fraction of true $b$-jets whose $b$-score exceeds the threshold, while the mis-ID rate is the fraction of non-$b$ jets passing the same threshold. In the efficiency range of $0.3$ to $0.8$, the JOI achieves misidentification rates that are one to three orders of magnitude lower than those of the multivariable analysis approach based on XGBoost refTDR_report.
@@
From the migration matrices at varying $R_{\mathrm{rad}}$ (see figure 4), we extract key flavour-tagging metrics as shown on the left-hand side of figure 5. As $R_{\mathrm{rad}}$ increases from $0.5$ to $2$, the $b\to c$, $g\to c$ and $c\to s$ misidentification rate rises from about $4\%$ to $7\%$, Meanwhile, the $c\to c$ tagging efficiency drops from $81\%$ to $70\%$, and $s\to s$ from $60\%$ to $54\%$. While the $g\to s$ rate remains stable around $14\%$. The trends indicate that reducing the vertex detector’s innermost radius significantly lowers the $b\to s$ and $c\to s$ misidentification rates, reflecting the crucial role of the vertex in distinguishing heavy-flavour jets from strange jets. In contrast, the impact of vertex detector optimisation on gluon-strange discrimination is limited as gluon jets predominantly originate from the primary vertex, and vertex detector upgrades have little effect on their identification.
@@
To further probe the global performance of JOI, we characterize its behaviour via the trace of $M_{11}$, denoted as $\mathrm{Tr}(M)$. This trace, which quantifies the total correct classification probability across all quark flavour categories, is analysed as a function of the geometry scaling factors $\log_{2}(\text{new}/\text{baseline})$ for both $R_{\rm rad}$ and $R_{\rm res}$. As shown on the right-hand side of figure  5, a linear regression to the $\mathrm{Tr}(M)$ yields:
@@
3.3 Performance of higgs benchmarks: $\nu\bar{\nu}H$ channel, $H\to c\bar{c}$ and $H\to s\bar{s}$ measurements
@@
with $S_{f\bar{f}}\in[0,1]$ and $\sum_{f}S_{f\bar{f}}=1$. For a given physics measurement, the candidate $H\to f\bar{f}$ signal events are defined as those satisfying $S_{f\bar{f}}$ over the threshold, where the threshold score is scanned to optimise the product $\epsilon\times P$, to balance the signal retention and background suppression, and thus the optimal measurement sensitivity. The efficiency $\epsilon$ is defined as the fraction of true $H\to f\bar{f}$ signal events whose $S_{f\bar{f}}$ value lies above the chosen threshold relative to the total number of simulated signal events. The purity $P$ is the fraction of selected events that truly originate from the $H\to f\bar{f}$ decay, i.e. the ratio of correctly identified signal events to all selected events including background contributions.
@@ 4 Summary and Discussion
Our study is based on fast simulations and currently considers only the major backgrounds from other Higgs decay modes, representing an optimistic yet well-motivated setup. Comparisons with full (Geant4-based) simulations confirm the reliability of our fast-simulation approach, with discrepancies at the 5% level. In a realistic environment, beam-induced and other Standard Model backgrounds would further dilute the sensitivity, and their quantitative impact, particularly the contributions related to the MDI design, still requires dedicated study. Consequently, the present results should be interpreted as upper-limit estimates achievable under idealised background conditions.
@@ Appendix A Comparison between Delphes and Geant4
Appendix B Simulation Study of the Beam-Induced Background Impact
The influence of beam-induced background on tracking performance has been qualitatively evaluated in the CEPC TDR CEPC2025TDR. As shown in Figure 10, a direct comparison of the impact-parameter resolutions ($d_{0}$ and $z_{0}$) with and without beam-induced background overlay demonstrates no noticeable degradation across the momentum range. Therefore, the pixel size and spatial resolution variations studied in this work remain valid, and the overall conclusions on vertex detector optimization are not affected by beam background effects.